OpenStudy (anonymous):

Can anyone solve it Laplace L(3t^2+3t^3+e^t+sin3t)

OpenStudy (unklerhaukus):

\[\mathcal L\{3t^2+3t^3+e^t+\sin3t\}\]\[\qquad=3\mathcal L\{t^2\}+3\mathcal L\{t^3\}+\mathcal L\{e^t\}+\mathcal L\{\sin 3t\}\]

OpenStudy (unklerhaukus):

\[\boxed{\mathcal L\big\{t^n\big\}=\dfrac{\Gamma(n+1)}{s^{n+1}}}\qquad\boxed{\mathcal L\big\{e^{-at}\big\}=\dfrac{1}{s+a}}\qquad\boxed{\mathcal L \big\{\sin(bt) \big\}=\dfrac{b}{s^2+b^2}}\]

OpenStudy (anonymous):

thank you so much dear..

OpenStudy (unklerhaukus):

\[\color{teal} {\ddot\smile}\]

OpenStudy (anonymous):

can i ask one more question..?

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

ok thnx.. d^3y/dt^3 + d^2y/dt^2 + dy/dt = sint

OpenStudy (unklerhaukus):

the laplace of?

OpenStudy (anonymous):

yup..

OpenStudy (unklerhaukus):

use \[\boxed{\mathcal L\big\{f'(t)\big\}=sF(s)-f(0)}\]\[\boxed{\mathcal L\big\{f^n(t)\big\}=s^nF(s)-s^{n-1}f(0)-s^{n-2}f'(0)\dots-sf^{n-2}(0)-f^{n-1}(0)}\]

hartnn (hartnn):

since initial conditions are not given, assume them to be 0.

OpenStudy (unklerhaukus):

hmm, the question should state the initial conditions

hartnn (hartnn):

which gives you, \(\boxed{\mathcal L\big\{f^n(t)\big\}=s^nF(s)}\)

hartnn (hartnn):

when not given, we can safely assume then to be 0.

OpenStudy (unklerhaukus):

i wouldn't assume that, i would have initial conditions in my final result

OpenStudy (abb0t):

hmmm....verrrry interesting symbols ya'll got there.

OpenStudy (anonymous):

i dont understand :(

OpenStudy (anonymous):

do u have the initial conditions?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

please give them

OpenStudy (anonymous):

Laplace d^3y/dt^3 + d^2y/dt^2 + dy/dt = sint

OpenStudy (anonymous):

there shud be more... r u give what y(0) is?

hartnn (hartnn):

dy/dt = f'(t) d^2y/dt^2 = f'' (t) d^3y/dt^3 = f'''(t) and then use, \(\boxed{\mathcal L\{f^n(t)\big\}=s^nF(s)-s^{n-1}f(0)-s^{n-2}f'(0)\dots-sf^{n-2}(0)-f^{n-1}(0)}\)

OpenStudy (anonymous):

given*

OpenStudy (anonymous):

DOnt know y(0) :(

OpenStudy (anonymous):

oh thnx hartnn

OpenStudy (anonymous):

i want thats,,step by step so thnx

OpenStudy (anonymous):

thanks to all who try to help me..

OpenStudy (anonymous):

i hope u can solve it now.... this might help, check out example 3: http://www2.fiu.edu/~aladrog/LaplaceTransDifferentialEq.pdf

OpenStudy (anonymous):

OK...thnx :)